OpenStudy (anonymous):

Solve the inequality 2(x + 6) greater than or equal to 3x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol.

OpenStudy (anonymous):

show the choices

OpenStudy (anonymous):

its a essay

OpenStudy (anonymous):

2(x+6)>=3x-2 3x-2x<=14 x<=14

OpenStudy (nurali):

2(x + 6) ≥ 3x - 2 2x+12≥3x-2 2x-3x≥-2-12 -1x≥-14 Divide both sides by -1 to get x alone x≤ 14

OpenStudy (anonymous):

2(x+6)>=3x-2 2x+12>=3x-2 3x-2x<=12+2 x<=14 close circle on 14:shading to the left

OpenStudy (anonymous):

@Game&twerk